OpenStudy (mackenzie6302):

evaluate the function h(x)=16-x, when x=-4 -64 12 20 64

OpenStudy (juscallmesteve):

20

OpenStudy (juscallmesteve):

h(x) = 16-(-4) h(x) = 16+4 h(x) = 20

OpenStudy (mackenzie6302):

thx